# Some insights into depth estimators for location and scatter in the multivariate setting

### v1

@@ 2 The location-scale model, depth and breakdown point
With this definition, we get very well known functionals for location and scale, as it is stated in the following lemma.
@@
If $\Theta$ denotes the set in which the parameters are assumed to lie in the central model (for instance, $\Theta=\mathbb{R}\times\left(0,\infty\right)$ in the location-scale model), the breakdown point of an estimator is a value that quantifies the level of contamination required to cause the estimator to move outside any compact subset of $\Theta$. A usual way to model the presence of spurious data in a sample is the $\varepsilon-$contamination neighborhood, which is defined to be,
$\mathcal{P}_{\varepsilon}\left(F_{0}\right)=\left\{\left(1-\varepsilon\right)F_{0}\left(\frac{\cdot-\mu_{0}}{\sigma_{0}}\right)+\varepsilon G\left(\cdot\right),\text{ }G\text{ any distribution on }\mathbb{R}\right\}.$
@@
$\varepsilon_{L}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:\sup_{\mathcal{P}_{\varepsilon}\left(F_{0}\right)}|\hat{\mu}\left(\cdot\right)|=\infty\right\},\text{ }$
$\varepsilon_{S}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:\sup_{\mathcal{P}_{\varepsilon}\left(F_{0}\right)}\max\left\{\hat{\sigma}\left(\cdot\right),\hat{\sigma}^{-1}\left(\cdot\right)\right\}=\infty\right\}.$
$\hat{\mu}_{1}$ as well as $\hat{\sigma}_{1}$ are estimators whose breakdown point is $0.5,$ the best breakdown point attainable for location and scale equivariant estimators. Let us suppose now that we bind together the estimation of both parameters in one expression, by considering the following second proposal for location-scale depth,
@@
$\left(\hat{\mu}_{2},\hat{\sigma}_{2}\right)=\arg\max_{\mu\in\mathbb{R},\sigma>0}D_{LS}^{2}\left(\mu,\sigma,P\right).$
It is easy to see that
@@
If $P=F_{0}\left(\sigma_{0}^{-1}\left(\cdot-\mu_{0}\right)\right),\left(\hat{\mu}_{2},\hat{\sigma}_{2}\right)=\left(\mu_{0},\sigma_{0}\right).$ If $P$ is a symmetric distribution, $\hat{\mu}^{2}=med\left(P\right),$ $\hat{\sigma}^{2}=med\left(P\right)-q_{1}\left(P\right),$ $q_{1}\left(P\right)$ the first quartile of $P.$ However, we can see in the Appendix that $\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)$ has breakdown point much lower than that of the $D_{LS}^{1}.$ We will assume that $F_{0}=\Phi$, the standard normal distribution. Let $\varepsilon_{0}$ be the solution to the equation
@@
with $h\left(x,y\right)=\Phi\left(x\right)-\Phi\left(\frac{y+x}{2}\right)$ and $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$
Theorem 1
The breakdown point $\varepsilon^{\ast}$ of $\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)$ is $\varepsilon^{\ast}=\varepsilon_{0}$ with $41/5<\varepsilon_{0}<1/$ 4 .
@@
The second proposal for depth estimators in the location-scale model yields estimators with much lower breakdown point compared to that of the first proposal. This phenomenum can be considered in a broader scenario showing a behavior also described in some other simultaneous procedures in the robust literature. For instance, a similar behavior is described in the simultaneous M-estimation for location and scale (Huber and Ronchetti, 2009, p.141) and in the context of simultaneous M-estimation for regression and scale in Maronna and Yohai (1991) which seems to prevent from doing simultaneous estimation in order to improve its robustness.
3 Maximum bias and concentration inequalities
The asymptotic breakdown point is an aspect of a more powerful tool to measure the performance of an estimator. The so-called maximum asymptotic bias function depicts more accurately the behavior of the estimator since it captures the global behavior by quantifying how much the estimator deviates in the whole contamination neighborhood under diferent levels of contaminations from the parameters at the central model. In spite of the deep understanding that the maxbias function provides, it is usually neglected because of the technicalities its derivation requires. .
$\mathcal{P}_{\varepsilon}\left(P_{0}\right)=\left\{\left(1-\varepsilon\right)P_{\mathbf{0}}\left(\mathbf{V}_{0}^{-1}\left(\cdot-\mathbf{\mu}_{0}\right)\right)+\varepsilon G\left(\cdot\right),\text{ }G\text{ any distribution on }\mathbb{R}^{p}\right\}$
@@ 3 Maximum bias and concentration inequalities
$\hat{\Gamma}\left(\mathcal{L}\left(A\mathbf{x+b}\right)\right)$ $=$ $A\hat{\Gamma}\left(\mathcal{L}\left(\mathbf{x}\right)\right)A^{t}$
We say that the functionals $\mathbf{\hat{\mu}}$ and $\hat{\Gamma}$ are Fisher consistent if
$\mathbf{\hat{\mu}}\left(P_{0}^{E}\right)=\mathbf{\mu}_{0}\text{ and }\hat{\Gamma}\left(P_{0}^{E}\right)=c\left(\mathbf{V}_{0}\mathbf{V}_{0}^{t}\right),\text{ }c>0\text{ for any }\mathbf{\mu}_{0},\mathbf{V}_{0}\in\mathbb{R}^{p\times p}\text{ invertible. }$.
The effect of the distorsion caused by having $P\in\mathcal{P}_{\varepsilon}\left(P_{0}^{E}\right)$, it is customary measured in the following invariant manner,
$b_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P\right)$ $=$ $\left(\mathbf{\hat{\mu}}\left(P\right)\mathbf{-\hat{\mu}}\left(P_{0}^{E}\right)\right)^{t}\left(\hat{\Gamma}\left(P_{0}^{E}\right)\right)^{-1}\left(\mathbf{\hat{\mu}}\left(P\right)\mathbf{-\hat{\mu}}\left(P_{0}^{E}\right)\right)$
$b_{S}\left(\hat{\Gamma},\varepsilon,P\right)$ $=$ $\sup_{\mathbf{u}\in\mathcal{S}^{p-1}}\frac{\mathbf{u}^{t}\hat{\Gamma}\left(P\right)\mathbf{u}}{\mathbf{u}^{t}\hat{\Gamma}\left(P_{0}\right)\mathbf{u}}$
$=$ $\lambda_{(1)}\left(\hat{\Gamma}\left(P_{0}\right)^{-1/2}\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1/2}\right)=\lambda_{(1)}\left(\hat{\Gamma}\left(P\right)\hat{\Gamma}\left(P_{0}\right)^{-1}\right),$
Then we can define the asymptotic biases for the largest and smallest eigenvalues as
$B_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P_{0}^{E}\right)$ $=$ $\sup_{P\in\mathcal{P}_{\varepsilon}}b_{L}\left(\mathbf{\hat{\mu}},\varepsilon,P\right)$
$B_{S}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $\sup_{P\in\mathcal{P}_{\varepsilon}}b_{S}\left(\hat{\Gamma},\varepsilon,P\right),\text{ }B_{I}\left(\hat{\Gamma},\varepsilon,,P_{0}^{E}\right)=\sup_{P\in\mathcal{P}_{\varepsilon}}b_{I}\left(\hat{\Gamma},\varepsilon,P\right).$
$B\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $\max\left\{B_{S}\left(\hat{\Gamma},\varepsilon\right),B_{I}\left(\hat{\Gamma},\varepsilon\right)\right\}$
$\varepsilon_{L}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:B_{L}\left(\hat{\Gamma},\varepsilon\right)=\infty\right\}$
$\varepsilon_{S}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:B_{S}\left(\hat{\Gamma},\varepsilon\right)=\infty\right\},\text{ }\varepsilon_{I}^{\ast}=\inf\left\{\varepsilon>0:B_{I}\left(\hat{\Gamma},\varepsilon\right)=\infty\right\},$
$\varepsilon^{\ast}$ $=$ $\min\left(\varepsilon_{S}^{\ast},\varepsilon_{I}^{\ast}\right).$
$B_{S}\left(\hat{\Gamma},\varepsilon,P_{0}^{E}\right)$ $=$ $B_{S}\left(\hat{\Gamma},\varepsilon,P_{0}\right)$
$B_{I}\left(\hat{\Gamma},\varepsilon,,P_{0}^{E}\right)$ $=$ $B_{I}\left(\hat{\Gamma},\varepsilon,,P_{0}\right)$
which entails that the maximum bias can be computed using $\mathbf{\mu}_{0}=\mathbf{0}$ and $\Sigma_{0}=I.$ The maximum bias for Tukey’s median turns out to be $B\left(\hat{\theta}_{T},\varepsilon,\Phi\right)=\Phi^{-1}\left(\frac{1+\varepsilon}{2\left(1-\varepsilon\right)}\right)$ (Chen and Tyler (2002)). Interest in scatter depth has been revitalized over the past decade, particularly following the seminal papers by Chen et al. (2018) and Paindaveine and Van Bever (2018), which explored different aspects of the concept of depth for multivariate scatter. If $X\sim P,$ $X\in\mathbb{R}^{p},\Gamma\in\mathcal{E}$, the depth of $\Gamma$ is taken to be
$D\left(\Gamma,P\right)=\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{\begin{array}[]{c}P\left(\left|\mathbf{u}^{t}\left(X-\mathbf{v}_{0}\left(P\right)\right)\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),\\ P\left(\left|\mathbf{u}^{t}\left(X-\mathbf{v}_{0}\left(P\right)\right)\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\end{array}\right\},$
with $\mathbf{v}_{0}$ a preliminary affine equivariant location functional to yield an affine equivariant multivariate scatter functional If location is known we can assume without loss of generality that $\mathbf{v}_{0}=\mathbf{0.}$ For the known location case, the depth and the deepest estimator are defined to be,
$D\left(\Gamma,P\right)$ $=$ $\inf_{\mathbf{u\in}\mathcal{S}^{p-1}}\min\left\{P\left(\left|\mathbf{u}^{t}X\right|^{2}\leq\mathbf{u}^{t}\Gamma\mathbf{u}\right),P\left(\left|\mathbf{u}^{t}X\right|^{2}\geq\mathbf{u}^{t}\Gamma\mathbf{u}\right)\right\}.$
$\hat{\Gamma}$ $=$ $\arg\underset{\Gamma\succeq\mathbf{0}}{\max}D\left(\Gamma,P\right).\text{ \ \ }D_{M}\left(P\right)=D\left(\hat{\Gamma},P\right),$
($\succeq$ defines the usual partial order in the set $\mathcal{E}$ of symmetric positive matrices).
Chen et al.. (2018) broke new ground by introducing a unified way to study the statistical convergence rate and robustness jointly. Given $\delta\in(0,1/2)$ let $\alpha=1-2\delta$ and $\varepsilon\in[0,1/2)$. Let $\mathcal{P}_{\varepsilon}\left(P_{0}\right)$ be the $\varepsilon-$contamination neighborhood with $P_{0}=N(\boldsymbol{\theta},\Sigma)$. Take $\mathcal{F}\left(M\right)$ as the set of symmetric and definite positive matrices $\Sigma$ such that the largest eigenvalue $\lambda_{1}(\Sigma)$ is less than a constant $M>0.$ They derived that, for $\varepsilon\in\left[0,\varepsilon^{\prime}\right],$ $\varepsilon^{\prime}<1/3$ and $\left(p+\log\left(1/\delta\right)\right)/n$ sufficiently small, there exists a constant $C>0$ (depending on $\varepsilon^{\prime}$ but independent of $p,n,\varepsilon),$ such that

### v2

@@ 5.1 Depth in the location-scale model
With this definition of depth, we obtain very well known functionals for location and scale, as it is stated in the following lemma.
Lemma 5.2
If $Y\sim P,$ then $\hat{\mu}_{1}=med_{P}\left(Y\right)$ and $\hat{\sigma}_{1}=med_{P}\left(\left|Y-med_{P}\left(Y\right)\right|\right)$ and $D_{LS}^{1}\left(\hat{\mu}_{1},\hat{\sigma}_{1},P\right)\geq 0.5.$
@@
A usual way to model the presence of spurious data in a sample is the $\varepsilon$-contamination neighborhood, which is defined to be,
@@
[…]
$\varepsilon_{S}^{\ast}$ $=$ $\inf\left\{\varepsilon>0:\sup_{\mathcal{P}_{\varepsilon}\left(F_{0}\right)}\max\left\{\hat{\sigma}\left(\cdot\right),\hat{\sigma}^{-1}\left(\cdot\right)\right\}=\infty\right\},\text{ }$
[…]
It is well known that $\hat{\mu}_{1}$ as well as $\hat{\sigma}_{1}$ are estimators whose breakdown point is $0.5,$ the best breakdown point attainable for location and scale equivariant estimators. We want to introduce now a slight modification which yields a substantial modification of the estimators. Let us suppose that we bind together the estimation of both parameters in one expression, by considering the following second proposal for location-scale depth,
@@
$\left(\hat{\mu}_{2},\hat{\sigma}_{2}\right)=\arg\max_{\mu\in\mathbb{R},\sigma>0}D_{LS}^{2}\left(\mu,\sigma,P\right),$
which we will call the modified location-scale deepest estimator. It is easy to see that
@@
If $P=F_{0}\left(\sigma_{0}^{-1}\left(\cdot-\mu_{0}\right)\right),$ then the modified estimator verifies that $\left(\hat{\mu}_{2}(P),\hat{\sigma}_{2}(P)\right)=\left(\mu_{0},\sigma_{0}\right)$ (that is, it is Fisher-consistent). If $P$ is a symmetric distribution, $\hat{\mu}^{2}=med\left(P\right),$ $\hat{\sigma}^{2}=med\left(P\right)-q_{1}\left(P\right),$ $q_{1}\left(P\right)$ is the first quartile of $P.$ However, we can prove that $\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)$ has breakdown point much lower than that of the $D_{LS}^{1}.$ We will assume that $F_{0}=\Phi$, the standard normal distribution. Let $\varepsilon_{0}$ be the solution to the equation
@@
with $h\left(x,y\right)=\Phi\left(x\right)-\Phi\left(\frac{y+x}{2}\right)$ and $y_{\delta}=\Phi^{-1}\left(\frac{\delta}{1-\delta}\right)$ a normal quantile, with $\delta<1/3.$ Then we obtain that the breakdown point for the modified estimator goes down.
Theorem 5.3
$\varepsilon^{\ast}\left(\hat{\mu}^{2},\hat{\sigma}^{2}\right)=\varepsilon_{0}$ with $1/5<\varepsilon_{0}<1/4.$
@@
The second proposal for depth estimators in the location-scale model yields estimators with much lower breakdown point compared to that of the first proposal. This phenomenon can be considered in a broader scenario showing a behavior also described in some other simultaneous procedures in the robust literature. While the estimation of $\mu$ and $\sigma$ is performed separately in (18), the estimation of both parameters is jointly performed in (5.1). For instance, a similar behavior is observed in the simultaneous M-estimation for location and scale (Huber and Ronchetti (2009), p.141) and in the context of simultaneous M-estimation for regression and scale in Maronna and Yohai (1991) which seems to prevent one from doing simultaneous estimation in order to improve its robustness.
6 Numerical study
In the previous section we derived the maximum bias curve for the deepest scatter matrix as well as its breakdown point. Only a few maxbias curves have been derived for multivariate dispersion measures, despite their being a very informative overall measure of robustness. Therefore, we cannot rely on comparisons among many theoretical maxbias curves to understand the behaviour of dispersion estimators.
On the one hand, maxbias curves are population measures; that is, they describe the behaviour of estimators as functionals over contamination neighborhoods. On the other hand, finite-sample effects should also be taken into account. In this section, we report the results of a Monte Carlo simulation study to investigate these effects, explicitly including the estimation of the unknown location parameter, whose impact is typically avoided in the maxbias derivation for scatter matrices because of the intractability to perform a theoretical analysis.
In Maronna and Yohai (2017) and Maronna et al. (2018), the behavior of scatter matrix estimators is studied under the assumption of known location, using the Kullback-Leibler divergence as a performance measure. Chen, Gao and Ren (2018a) addresses the performance of estimators under unknown location by using the operator norm. We consider the empirical version of the bias given in formulae (3.1) and (3.1) as performance measures. Following Maronna and Yohai (1995), we also use the condition number of an estimator as a measure of asymptotic bias.
In the following subsections, we describe the estimators and contamination scenarios, explain how the estimated bias measures are empirically computed from replicated samples of the contaminated model, and then report and discuss the main results.
6.1 The estimators
Let us consider $\mathbf{X}=(X_{1},\ldots,X_{p})^{T}$ a random vector and $\mathbf{X}_{1},\ldots,\mathbf{X}_{n}$ a random sample from $\mathbf{X}$. Let $\mathbb{X}=\{\mathbf{x}_{1},\ldots,\mathbf{x}_{n}\}\subset\mathbb{R}^{p}$ denote a dataset. Consider $\mathbf{x},\boldsymbol{\mu}\in\mathbb{R}^{p}$ and $\Sigma\in\mathbb{R}^{p\times p}$, and let $d(\mathbf{x},\boldsymbol{\mu},\Sigma)=(\mathbf{x}-\boldsymbol{\mu})^{\prime}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})$ be the squared Mahalanobis distance. Let $d_{i}=d(\mathbf{x}_{i},\boldsymbol{\mu},\Sigma)$ and $\mathbf{d}(\boldsymbol{\mu},\Sigma)=(d_{1},\ldots,d_{n}),$ $i=1,\dots,n$.
We include several estimators in our simulation study. To compute them, we use functions from R packages available on the Comprehensive R Archive Network (CRAN), with default argument values in all cases. A brief description of them is as follows.
  • 1.
    The sample covariance matrix (SCOV), defined by $S_{n}=\frac{1}{n}\sum_{i=1}^{n}(\boldsymbol{x}_{i}-\bar{\mathbf{x}}_{n})(\mathbf{x}_{i}-\bar{\mathbf{x}}_{n})^{\prime}$, where $\bar{\mathbf{x}}_{n}$ is the sample mean, included to provide a benchmark for comparison with the robust estimators that follow.
  • 2.
@@ 6.1 The estimators
    $\left(\hat{\boldsymbol{\mu}},\hat{\Sigma}\right)=\arg\underset{\left(\boldsymbol{t},\Sigma\right)\in C_{h,c}}{\min}\det\left(\Sigma\right).$
    $c$ is chosen to yield a consistent estimator $\hat{\Sigma}$ at the central model, and $h=[(n+p+1)/2]$ ensures the maximal breakdown point, as described in Van Aelst and Rousseeuw (2009). MVE is computed in R using the function fastmve from the package RobStatTM (Yohai et al., 2024).
  • 3.
    The minimum covariance determinant estimator (MCD), was proposed by Rousseeuw (1985)). Given the $n$ observations, this method finds $h$ observations $\mathbf{x}_{i_{1}},\ldots,\mathbf{x}_{i_{h}}$ with sample covariance matrix $S_{h}$ having the lowest determinant, with $\lfloor(n+p+1)/2\rfloor\leq h\leq n.$ $S_{h}$ and the sample mean $\bar{\mathbf{x}}_{n}$ of the $h$ observations are the MCD estimators for multivariate scatter and location. It is computed in R using the CovMcd function from the package rrcov developed by Todorov and Filzmoser (2009).
  • 4.
    S-estimators (SE) were introduced by Davies (1987). Given the constant $\delta\in(0,1),$ an M-scale $S=S(\mathbf{d}(\boldsymbol{\mu},\Sigma))$ is defined through the equation
    $\frac{1}{n}\sum_{i=1}^{n}\rho\!\left(\frac{d_{i}}{S}\right)=\delta,$
[…]
    $(\widehat{\boldsymbol{\mu}},\widetilde{\Sigma})=\arg\min_{\boldsymbol{\mu}\in\mathbb{R}^{p},\Sigma\in\mathcal{E},|\Sigma|=1}S\bigl(\mathbf{d}(\boldsymbol{\mu},\Sigma)\bigr),$
    where $\delta$ controls the breakdown point, then it is taken as $\delta=1/2$. Hence, the S-estimator for multivariate scatter $\Sigma$ is defined as $\widehat{\Sigma}=S\,\widetilde{\Sigma}$. SE is computed using the function CovSest from the package rrcov, choosing method = "bisquare", with the bisquare function.
    $\rho(t)=\begin{cases}1-(1-t)^{3}&\text{if }t\leq 1,\\ 1&\text{if }t>1.\end{cases}$
  • 5.
    S-estimators with non-monotonic weight functions (Rocke ) were introduced by Rocke (1996). Note that S-estimators satisfy the system
    $\frac{1}{n}\sum_{i=1}^{n}W\left(\frac{d_{i}}{S}\right)(\mathbf{x}_{i}-\boldsymbol{\mu})(\mathbf{x}_{i}-\boldsymbol{\mu})^{\prime}$ $=$ $\Sigma,$
    $\frac{1}{n}\sum_{i=1}^{n}W\left(\frac{d_{i}}{S}\right)(\mathbf{x}_{i}-\boldsymbol{\mu})$ $=$ $0,$
    where $W=\rho^{\prime}$ is usually called a weight function. Rocke is computed using the function CovSest from the package rrcov with option method ="rocke", that by default sets $\alpha=0.1$. This routine considers a modification of Rocke’s “biflat” weight function given by
    $W(t)=\left[1-\left(\frac{t-1}{\gamma}\right)^{2}\right]\mathrm{1}_{(1-\gamma,1+\gamma)}(t),$
    with tuning constant $\gamma=\min\!\left(1,\frac{\chi_{p,1-\alpha}^{2}}{p}-1\right)$ where $\alpha$ is chosen to control efficiency.
  • 6.
    MM-estimators (MM) were proposed by Lopuhaa (1992). Let $\left(\widehat{\boldsymbol{\mu}}_{0},\widehat{\Sigma}_{0}\right)$ be an initial high breakdown point estimator and the squared Mahalanobis distances $d_{i}^{0}=d\left(\mathbf{x}_{i},\widehat{\boldsymbol{\mu}}_{0},\widehat{\Sigma}_{0}\right)$. Consider an M-scale $S^{0}$ solving
    $\frac{1}{n}\sum_{i=1}^{n}\rho\left(\frac{d_{i}^{0}}{S}\right)=\delta.$
    In the end, consider the location and shape estimator defined as